Maryann is planning a wedding anniversary gift of a trip to Hawaii for her husband at the end of 5 years. She will have enough to pay for the trip if she invests $5,000 per year until that anniversary and plans to make her first $5,000 investment on their first anniversary. Assume her investment earns a 4 percent interest rate, how much will she have saved for their trip if the interest is compounded in each of the following ways?

a. Annually b. Quarterly c. Monthly
1
Expert's answer
2020-10-27T08:09:12-0400

Assuming that Maryann invests her $5,000 at the beginning of each year, and for 5 years. The investment is therefore an annuity due, and we need to find future value at the end of 5 years. For each year, we are provided:

PMT = $5,000

i = 4%

n = 5 years.


The future value "(FV)" of an annuity due is found by the formula:

"FV = PMT \left[ \dfrac {\left(1+ \dfrac {i}{m}\right)^{mn}-1}{\left(\dfrac {i}{m}\right)} \right]\left(1 + \dfrac {i}{m}\right)"

Where, "i" is the annual interest rate, "n" is the number of years, "PMT" is the annual payment, and "m" is the number of compounding periods.


"\bold {(a) \space Annual \space Compounding}"

"\bold {Answer}"

"FV =\$28,164.88"


"\bold {Solution}"

"m = 1"


"FV = 5,000 \left[\dfrac {(1+0.04)^5- 1}{0.04}\right](1+0.04)"

"FV = 5,000 \left[\dfrac {1.04^5- 1}{0.04}\right](1.04)"

"= \$28,164.877312"

"=\bold {\$28,164.88}"



"\bold {(b) \space Quarterly \space Compounding}"


"\bold {Answer}"

"FV = \$111,195.97"


"\bold {Solution}"

"m = 4"


"FV = 5,000 \left[ \dfrac {\left(1+ \dfrac {0.04}{4}\right)^{(5×4)}-1}{\left(\dfrac {0.04}{4}\right)} \right]\left(1 + \dfrac {0.04}{4}\right)"

"FV = 5,000 \left[\dfrac {(1+0.01)^{20} - 1}{0.01}\right](1+0.01)"


"FV = 5,000 \left[\dfrac {1.01^{20}- 1}{0.01}\right](1.01)"


"= \$111,195.97015"

"= \bold {\$111,195.97}"



"\bold {(c) \space Monthly \space Compounding}"


"\bold {Answer}"

"FV = \$332,599.87"


"\bold {Solution}"

"m = 12"


"FV = 5,000 \left[ \dfrac {\left(1+ \dfrac {0.04}{12}\right)^{(5×12)}-1}{\left(\dfrac {0.04}{12}\right)} \right]\left(1 + \dfrac {0.04}{12}\right)"


"FV = 5,000 \left[ \dfrac {\left(1+ \dfrac {0.04}{12}\right)^{60}-1}{\left(\dfrac {0.04}{12}\right)} \right]\left(1 + \dfrac {0.04}{12}\right)"

"=\$332,599.873883"

"= \bold {\$322,599.87}"